Find the x-intercept of -6x + 9y = -18 Write answer as (x,y)
Answer by math_tutor2020(3817)


The x-intercept always occurs when y = 0.
This is where the curve either touches the x axis, or crosses over the x axis.

-6x+9y = -18
-6x+9*0 = -18
-6x = -18
x = -18/(-6)
x = 3
The x-intercept is located at (3,0)

Below is a graph to verify the answer.
Note that -6x+9y = -18 solves to y = (2/3)x-2

GeoGebra and Desmos are two of many options if you need a graphing calculator.
Or you can go old-school with something like a TI83.

--------------------------------------------------------------------------

Answer: (3,0)